# A Survey of Mamba

### v3

@@ 5.2.2. Graph-structured Data
Graph modeling has found extensive utility in managing complex structures and relationships, including applications in domains like social networks (Fan et al., 2019b, 2020), recommender systems (Fan et al., 2022), and molecular interactions (Huang et al., 2020). Due to the powerful capabilities of Mamba in long-range modeling and high efficiency, several pioneering investigations have embraced the selective State Space Model (SSM) for non-sequential graph data (Liu et al., 2024a). These studies utilize state space models to encode context through hidden states during recurrent scans, allowing for input flow control, which resembles attention sparsification on graphs, presenting a data-dependent node selection process within graph modeling contexts (Ye and Ji, 2021). Moreover, Mamba is anticipated to enhance model efficiency during large-graph training tasks. For example, Graph-Mamba (Wang et al., 2024e) introduces a novel Mamba-based block as a foundational component for graph modeling. This block combines a graph flattening mechanism with the selection mechanism offered by Mamba, transforming sub-graphs into node sequences and facilitating input-dependent context filtering, respectively. In a recent work, Behrouz and Hashemi (2024) propose a Graph Mamba Network (GMN), a new graph neural network format based on selective SSMs. The authors reformulate the selective SSM into a graph learning format and provide theoretical justification for the power of the proposed network. By addressing the emerging challenges in crucial steps of graph message passing, GMNs achieve remarkable performance in various aspects, surpassing GNNs and Transformer-based models in multiple benchmark datasets with diverse graph scales. Furthermore, Huang et al. (2024a) introduce the Graph State Space Convolution (GSSC) as a systematic extension of SSMs tailored for graph-structured data. Specifically, GSSC incorporates distance-based graph convolution kernels into the SSM cell, aiming at enhancing expressive power and capturing long-range dependencies. Through assessments conducted on ten benchmark datasets, the study (Huang et al., 2024a) underscores the potential of GSSC as a potent and scalable model for graph machine learning.
@@ 5.3. Multimodal Data
Integrating multiple modalities, such as language (sequential data) and images (non-sequential data), offers valuable and complementary information for artificial intelligence perception and scene understanding. Recently, there has been significant research attention on Multimodal Large Language Models (MLLMs) that inherit the advanced capabilities of Large Language Models (Wu et al., 2023), including powerful language expression and logical reasoning. While Transformers have been the dominant approach in this field, Mamba has emerged as a strong competitor by demonstrating impressive performance in aligning mixed-source data and achieving linear complexity scaling in sequence length, which makes Mamba a promising alternative to Transformers for multimodal learning (Yang et al., 2024c; Liu et al., 2024b). For example, Qiao et al. (2024) propose VL-Mamba to explore the utilization of Mamba’s efficient architecture for solving vision-language tasks, utilizing the pre-trained Mamba model for language understanding and incorporating a connector module to align visual patches with language tokens. Wang et al. (2024d) propose Text-controlled Motion Mamba (Wang et al., 2024d), which leverages Mamba to dynamically capture global temporal information based on text queries to enhance human motion understanding. Additionally, Fusion-Mamba (Dong et al., 2024) and Sigma (Wan et al., 2024) have tried to fuse complementary information from different modalities such as thermal, depth, and RGB. Fusion-Mamba focuses on improving object detection, while Sigma aims to enhance semantic segmentation.
@@ 6.4. Drug Discovery
Protein design, molecular design, and genomic analysis are pivotal in advancing drug discovery and biotechnology (Scott et al., 2016; Li et al., 2024d). Leveraging the Mamba-based model significantly reduces the complexities of modeling long sequences in these domains (Peng et al., 2024, 2024; Guo and Schwaller, 2024; Schiff et al., 2024, 2024). Specifically, PTM-Mamba (Peng et al., 2024) and ProtMamba (Sgarbossa et al., 2024) are protein language models based on the Mamba architecture. PTM-Mamba utilizes bidirectional gated Mamba blocks and structured state space models, efficiently processing long sequences while reducing computational demands. ProtMamba is designed to be homology-aware yet alignment-free, adept at handling extensive contexts across hundreds of protein sequences. Both models maintain high efficiency and accuracy even with large data sets, providing critical tools for protein design. Meanwhile, generative molecular design aims to simulate molecules with tailored property profiles from a specific distribution. However, current models lack the efficiency required to optimize high-fidelity oracles, directly resulting in low success rates. Saturn (Guo and Schwaller, 2024), applying the Mamba architecture, utilizes its linear complexity and computational efficiency to surpass 22 competing models in drug discovery. Furthermore, understanding genomes is crucial for gaining insights into cellular biology. Challenges in genomic modeling include capturing interactions between distant tokens, considering the impacts of both upstream and downstream regions, and ensuring the complementarity of DNA sequences. Caduceus (Schiff et al., 2024) and MSAMamba (Thoutam and Ellsworth, [n. d.]), both leveraging the Mamba model, excel in addressing these challenges. Caduceus, a DNA foundation model, enhances the Mamba architecture with BiMamba and MambaDNA components for bi-directional modeling and ensuring reverse complement equivariance, significantly outperforming existing models in long-range genomic tasks. Similarly, MSAMamba (Thoutam and Ellsworth, [n. d.]) addresses the limitations of transformer-based models for DNA multiple sequence alignments by implementing a selective scan operation along the sequence dimension. This design extends the training context length of previous methods by eightfold, allowing a more comprehensive analysis of extensive DNA sequences.
@@ 6.5. Recommender Systems
Recommender Systems are widely utilized in e-commerce (Zhang et al., 2024e; Zhou et al., 2018; Chen et al., 2023) and social networks (Fan et al., 2019c, a, 2018), aiming to capture users’ evolving preferences and the interdependencies among their past behaviors (Zhao et al., 2024b; Fan et al., 2022). Although transformer-based models have demonstrated effectiveness in recommender systems (Sun et al., 2019), they face computational efficiency challenges because of the quadratic complexity of attention mechanisms, especially when dealing with longer sequences of behaviors. Recently, several Mamba-based models have been applied to analyze long-term user behavior for personalized recommendations (Yang et al., 2024d; Liu et al., 2024a; Wang et al., 2024c; Su and Huang, 2024; Cao and Zhang, 2024). For instance, Mamba4Rec (Liu et al., 2024a) pioneers the use of selective state space models for efficient sequential recommendation, enhancing model performance while maintaining inference efficiency. Similarly, RecMamba (Yang et al., 2024d) explores Mamba’s effectiveness in lifelong sequential recommendation scenarios (i.e., sequence length $\geq$ 2k), achieving comparable performance to benchmark models while cutting down training time by 70% and reducing memory costs by 80%. Furthermore, EchoMamba4Rec (Wang et al., 2024c) integrates a bidirectional Mamba module with frequency-domain filtering to accurately capture intricate patterns and interdependencies within user interaction data. It demonstrates superior performance over existing models, delivering more precise and personalized recommendations. Additionally, Mamba4KT (Cao and Zhang, 2024) is designed explicitly for knowledge tracing in intelligent education, leveraging the Mamba model to capture enduring correlations between exercises and student knowledge levels. As educational datasets expand, this method suggests a promising avenue for enhancing prediction accuracy, model efficiency, and resource utilization in knowledge tracing research.

### v4

@@ 5.2.2. Graph-structured Data
Graph modeling has found extensive utility in managing complex structures and relationships, including applications in domains like social networks (Fan et al., 2019b, 2020), recommender systems (Fan et al., 2022), and molecular interactions (Huang et al., 2020). Due to the powerful capabilities of Mamba in long-range modeling and high efficiency, several pioneering investigations have embraced the selective State Space Model (SSM) for non-sequential graph data (Liu et al., 2024a). These studies utilize state space models to encode context through hidden states during recurrent scans, allowing for input flow control, which resembles attention sparsification on graphs, presenting a data-dependent node selection process within graph modeling contexts (Ye and Ji, 2021). Moreover, Mamba is anticipated to enhance model efficiency during large-graph training tasks. For example, Graph-Mamba (Wang et al., 2024g) introduces a novel Mamba-based block as a foundational component for graph modeling. This block combines a graph flattening mechanism with the selection mechanism offered by Mamba, transforming sub-graphs into node sequences and facilitating input-dependent context filtering, respectively. In a recent work, Behrouz and Hashemi (2024) propose a Graph Mamba Network (GMN), a new graph neural network format based on selective SSMs. The authors reformulate the selective SSM into a graph learning format and provide theoretical justification for the power of the proposed network. By addressing the emerging challenges in crucial steps of graph message passing, GMNs achieve remarkable performance in various aspects, surpassing GNNs and Transformer-based models in multiple benchmark datasets with diverse graph scales. Furthermore, Huang et al. (2024a) introduce the Graph State Space Convolution (GSSC) as a systematic extension of SSMs tailored for graph-structured data. Specifically, GSSC incorporates distance-based graph convolution kernels into the SSM cell, aiming at enhancing expressive power and capturing long-range dependencies. Through assessments conducted on ten benchmark datasets, the study (Huang et al., 2024a) underscores the potential of GSSC as a potent and scalable model for graph machine learning.
@@ 5.3. Multimodal Data
Integrating multiple modalities, such as language (sequential data) and images (non-sequential data), offers valuable and complementary information for artificial intelligence perception and scene understanding. Recently, there has been significant research attention on Multimodal Large Language Models (MLLMs) that inherit the advanced capabilities of Large Language Models (Wu et al., 2023), including powerful language expression and logical reasoning. While Transformers have been the dominant approach in this field, Mamba has emerged as a strong competitor by demonstrating impressive performance in aligning mixed-source data and achieving linear complexity scaling in sequence length, which makes Mamba a promising alternative to Transformers for multimodal learning (Yang et al., 2024c; Liu et al., 2024b). For example, Qiao et al. (2024) propose VL-Mamba to explore the utilization of Mamba’s efficient architecture for solving vision-language tasks, utilizing the pre-trained Mamba model for language understanding and incorporating a connector module to align visual patches with language tokens. Wang et al. (2024f) propose Text-controlled Motion Mamba (Wang et al., 2024f), which leverages Mamba to dynamically capture global temporal information based on text queries to enhance human motion understanding. Additionally, Fusion-Mamba (Dong et al., 2024) and Sigma (Wan et al., 2024) have tried to fuse complementary information from different modalities such as thermal, depth, and RGB. Fusion-Mamba focuses on improving object detection, while Sigma aims to enhance semantic segmentation.
@@ 6.4. Drug Discovery
Protein design, molecular design, and genomic analysis are pivotal in advancing drug discovery and biotechnology (Scott et al., 2016; Li et al., 2024d). Leveraging the Mamba-based model significantly reduces the complexities of modeling long sequences in these domains (Peng et al., 2024, 2024; Guo and Schwaller, 2024; Schiff et al., 2024, 2024). Specifically, PTM-Mamba (Peng et al., 2024) and ProtMamba (Sgarbossa et al., 2024) are protein language models based on the Mamba architecture. PTM-Mamba utilizes bidirectional gated Mamba blocks and structured state space models, efficiently processing long sequences while reducing computational demands. ProtMamba is designed to be homology-aware yet alignment-free, adept at handling extensive contexts across hundreds of protein sequences. Both models maintain high efficiency and accuracy even with large data sets, providing critical tools for protein design. Meanwhile, generative molecular design aims to simulate molecules with tailored property profiles from a specific distribution. However, current models lack the efficiency required to optimize high-fidelity oracles, directly resulting in low success rates. Saturn (Guo and Schwaller, 2024), applying the Mamba architecture, utilizes its linear complexity and computational efficiency to surpass 22 competing models in drug discovery. Furthermore, understanding genomes is crucial for gaining insights into cellular biology. Challenges in genomic modeling include capturing interactions between distant tokens, considering the impacts of both upstream and downstream regions, and ensuring the complementarity of DNA sequences. Caduceus (Schiff et al., 2024) and MSAMamba (Thoutam and Ellsworth, [n. d.]), both leveraging the Mamba model, excel in addressing these challenges. Caduceus, a DNA foundation model, enhances the Mamba architecture with BiMamba and MambaDNA components for bi-directional modeling and ensuring reverse complement equivariance, significantly outperforming existing models in long-range genomic tasks. Similarly, MSAMamba (Thoutam and Ellsworth, [n. d.]) addresses the limitations of transformer-based models for DNA multiple sequence alignments by implementing a selective scan operation along the sequence dimension. This design extends the training context length of previous methods by eightfold, allowing a more comprehensive analysis of extensive DNA sequences. Another illustration is SMILES-Mamba (Xu et al., 2024b), a system that pre-trains and fine-tunes Mamba models to forecast the absorption, distribution, metabolism, excretion, and toxicity characteristics of small-molecule drugs.
@@ 6.5. Recommender Systems
Recommender Systems are widely utilized in e-commerce (Zhang et al., 2024e; Zhou et al., 2018; Chen et al., 2023) and social networks (Fan et al., 2019c, a, 2018), aiming to capture users’ evolving preferences and the interdependencies among their past behaviors (Zhao et al., 2024b; Fan et al., 2022). Although transformer-based models have demonstrated effectiveness in recommender systems (Sun et al., 2019), they face computational efficiency challenges because of the quadratic complexity of attention mechanisms, especially when dealing with longer sequences of behaviors. Recently, several Mamba-based models have been applied to analyze long-term user behavior for personalized recommendations (Yang et al., 2024d; Liu et al., 2024a; Wang et al., 2024d; Su and Huang, 2024; Cao and Zhang, 2024). For example, Mamba4Rec (Liu et al., 2024a) pioneers the use of selective state space models for efficient sequential recommendation, enhancing model performance while maintaining inference efficiency. Similarly, RecMamba (Yang et al., 2024d) explores Mamba’s effectiveness in lifelong sequential recommendation scenarios (i.e., sequence length $\geq$ 2k), achieving comparable performance to benchmark models while cutting down training time by 70% and reducing memory costs by 80%. Furthermore, EchoMamba4Rec (Wang et al., 2024d) integrates a bidirectional Mamba module with frequency-domain filtering to accurately capture intricate patterns and interdependencies within user interaction data. It demonstrates superior performance over existing models, delivering more precise and personalized recommendations. Additionally, Mamba4KT (Cao and Zhang, 2024) is designed explicitly for knowledge tracing in intelligent education, leveraging the Mamba model to capture enduring correlations between exercises and student knowledge levels. As educational datasets expand, this method suggests a promising avenue for enhancing prediction accuracy, model efficiency, and resource utilization in knowledge tracing research. More recently, SSD4Rec (Qu et al., 2024c) proposes a novel pure-SSM structure for efficient sequential recommendation, which harnesses the advanced nature of Structured State-Space Duality to enable attention-like parallel computation on item sequences while scaling linearly with the input lengths. SSD4Rec allows variable-length input without any padding/truncation and achieves comprehensive long-range user behavior modeling from bi-directional perspectives.